The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: Eva K. Grebel, You-Hua Chu, and Dominik J. Bomans

                                 mcnews@astro.uiuc.edu
                      http://www.astro.uiuc.edu/mcnews/MCNews.html
                         http://www.astro.uni-bonn.de/~mcnews/

No. 17 February 8, 1998

Table of Contents

Abstracts of 10 refereed papers 1
Thesis abstract  8
Meeting announcement  9

Abstracts of Refereed Papers

Starburst-like Dust Extinction in the Small Magellanic Cloud

K.D. Gordon and G.C. Clayton

Department of Physics & Astronomy, Louisiana State University, Baton Rouge, LA 70803, USA

The recent discovery that the UV dust extinction in starburst galaxies is similar to that found in the Small Magellanic Cloud (SMC) motivated us to re-investigate the ultraviolet (UV) extinction found in the SMC. We have been able to improve significantly on previous studies by carefully choosing pairs of well matched reddened and unreddened stars. In addition, we benefited from the improved S/N of the NEWSIPS IUE data and the larger sample of SMC stars now available. Searching the IUE Final Archive, we found only four suitable early-type stars that were significantly reddened and had well matched comparison stars. The extinction for three of these stars is remarkably similar. The curves are roughly linear with λ^{−1} and have no measurable 2175 Å bump. The fourth star has an extinction curve with a significant 2175 Å bump and weaker far-UV extinction. The dust along all four sightlines is thought to be local to the SMC. There is no significant Galactic foreground component. The first three stars lie in the SMC Bar and the line-of-sight for each of them passes through regions of recent star formation. The fourth star belongs to the SMC Wing and its line-of-sight passes though a much more quiescent region. Thus, the behavior of the dust extinction in the SMC supports a dependence of dust properties on star formation activity. However, other environmental factors (such as galactic metallicity) must also be important. Dust in the 30 Dor region of the LMC, where much more active star formation is present, does not share the extreme extinction properties seen in SMC dust.

Accepted by: The Astrophysical Journal, (20 June 1998)

For preprints, contact gordon@fenway.phys.lsu.edu
Also available from the URL http://snake.phys.lsu.edu/∼gordon/papers/SMC ext.html
or by anonymous ftp at ftp://fenway.phys.lsu.edu/kdg/SMC ext.ps.gz

Extended Red Emission (ERE) detected in the 30 Doradus nebula

S. Darbon, J.-M. Perrin, and J.-P. Sivan

Observatoire de Haute Provence du CNRS, F-04870 Saint Michel l’Observatoire, France

We present low-dispersion visible spectra of two regions of the 30 Doradus nebula. When corrected for atomic continuum emission and divided by the spectrum of the exciting and illuminating stars, one of the two spectra clearly shows Extended Red Emission (ERE) superimposed on the scattering component. This ERE band peaks around 7270 Å and is 1140 Å wide. HAC grains are found to explain the observed spectra in terms of scattering and luminescence.

Accepted by: Astronomy & Astrophysics For preprints, contact sivan@obs-hp.fr Also available from the URL http://www.obs-hp.fr/

Ejected Nebulae as Probes of the Evolution of Massive Stars in the Large Magellanic Cloud

Linda J. Smith^{1}, Antonella Nota^{2}, Anna Pasquali^{2}, Claus Leitherer^{2}, Mark Clampin^{2} and Paul A. Crowther^{1}

1 Department of Physics and Astronomy, University College London, Gower St., London WC1E 6BT, UK 2 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA

We present new HST/FOS and ground-based long slit spectroscopic observations of the nebulae around the two LMC Luminous Blue Variables R127 (=HDE269858) and R143 (=HDE269929), and the Ofpe/WN9 star S119 (=HDE269687).

We have used the ground-based long slit spectra to investigate the kinematics of the nebulae around R127 and R143. We find that the circumstellar environment of R127 is complex and suggestive of two discrete ejections in the mass loss history of the central star. There is an inner expanding shell, with a radius of 0.6 pc, an expansion velocity of 29 km s^{−1} and a dynamical age of 2 × 10^{4} yr. There is also material beyond the bright, inner nebula which may represent an outer expanding shell with a radius of 1.9 pc, and a dynamical age of 7 × 10^{4} yr. For R143, we find that the velocity field in the northern and southern circumstellar regions (1–20′′) is predominantly constant. Therefore, the nebular emission previously believed to be associated with R143 (the fingers) is most likely part of the 30 Doradus complex, with the exception of a small emission line region located ≃ 2′′ north of the star. This compact region displays a blueshifted motion with a differential velocity of ≃ 130 km s^{−1} with respect to the central star. The spectrum of this clump shows a high [N II]/Hα ratio, suggestive of nitrogen-enriched material that has been ejected from the star some 3.7 × 10^{3} yrs ago.

We have used the optical HST+FOS spectra to determine reddenings, electron temperatures and densities, and N and O abundances for the three nebulae. For R143, our derived abundances indicate that the region we observed to the south of the star is 30 Doradus H II gas since it shows a typical LMC H II region N/O ratio. For R127, we find that N is enriched by a factor of 10.7 ± 2.2, and O depleted by a factor of 2.0 ± 1.0 with N/O= 0.89 ± 0.40. For the S119 nebula, we do not have a secure electron temperature but find N/O=1.41–2.45, similar to the value found for R127. We compare the nebular abundances with the expected surface abundances of LBVs and discuss the likely stellar evolutionary state at the time of nebular ejection. If the atmospheres of LBVs consist of CNO-processed material, this implies that the event that formed the nebula took place before, or at the very start of, the LBV phase.

Comparison of the observed nebular abundances with other objects shows that the abundance pattern for the R127 nebula is remarkably similar to that determined for the inner ring of SN 1987A (Fransson et al. 1989; Panagia et al. 1997) which is thought to be composed of RSG wind material. This coincidence, together with the observed low expansion velocity, suggests that the nebula was once the CN-processed convective envelope of an RSG. We consider whether the R127 nebular parameters can be reproduced with the evolutionary models of Meynet et al. (1994), and find that the pre-LBV mass loss has to be low enough for redward evolution such that a surface N/O ratio of ∼ 1 is achieved while the star is a cool supergiant inside the Humphreys-Davidson limit. Our findings are broadly in accord with the model of Stothers & Chin (1996) which incorporates a brief RSG phase. We also consider the possibility of a pseudo-RSG phase by discussing the suggestion of Owocki & Gayley (1997) that a star in a super-Eddington condition may develop a very extended outer convective envelope which may become detached.

We also present a coarse analysis of the FOS stellar spectrophotometry using the Q-method to estimate reddenings and spectral types, and compare our results with more sophisticated methods.

Submitted to: The Astrophysical Journal

Preprints from ljs@star.ucl.ac.uk
or by anonymous ftp to ftp.star.ucl.ac.uk, in pub/ljs/lmc/paper.ps.gz

The OB associations LH 101 and LH 104 in the HII region N158 of the LMC

G. Testor^{1} and V. Niemela^{2}

1 Observatoire de Paris, section de Meudon, F-92195 Meudon Cedex, DAEC, France 2 Facultad de Ciencias Astronómicas y Geofísicas, Universidad Nacional de La Plata, 1900 La Plata, Argentina

We present photometric and spectroscopic observations of stars in the Large Magellanic Cloud OB associations LH 101 and LH 104, located in the HII region N158, which we have also imaged. From our observations we have constructed upper H-R diagrams for these OB associations, which we find to consist mainly of three populations, one of 2-6 Myr for the stars inside the northern bubble (LH 104), and two populations in the southern HII region (LH 101), one of ≤ 2 Myr and the other one aged 3-6 Myr. We have obtained for LH 101 a normal IMF, with a slope of Γ = −1.29 ± 0.20 whereas for LH 104 the IMF is flatter with a slope of Γ = −1.05 ± 0.12. These IMF slopes are consistent with that of other OB associations in the LMC. Our observations reveal in the region of LH 101 the presence of both unevolved and evolved very massive stars, whose ionizing flux is in excess of that derived from our Hβ images of the HII region. The north-west nebulosity in the region of LH 101 thus appears to be matter bound.

Accepted by: Astronomy and Astrophysics (Supplement Series)

For preprints, contact gerard.testor@obspm.fr

Evolved Massive Stars in the Local Group:

I. Identification of Red Supergiants in NGC 6822, M31, and M33

Philip Massey

Kitt Peak National Observatory, NOAO, P.O. 26732, Tucson, AZ 85726-6732, USA

Knowledge of the red supergiant (RSG) population of nearby galaxies allows us to probe massive star evolution as a function of metallicity; however, contamination by foreground Galactic dwarfs dominates surveys for red stars in Local Group galaxies beyond the Magellanic Clouds. Model atmospheres predict that low-gravity supergiants will have B − V values that are redder by several tenths of a magnitude than foreground dwarfs at a given V − R color, a result that is largely independent of reddening. We conduct a BV R survey of several fields in the Local Group galaxies NGC 6822, M 33, and M 31, as well as neighboring control fields, and identify RSG candidates from CCD photometry. The survey is complete to V = 20.5, corresponding to M_{V} = −4.5 or an M_{bol} of −6.3 for the reddest stars. Followup spectroscopy at the Ca II triplet of 130 stars is used to demonstrate that our photometric criterion for identifying RSGs is highly successful (96% for stars brighter than V = 19.5; 82% for for V =19.5-20.5). Classification spectra are also obtained for a number of stars in order to calibrate color with spectral type empirically. We find that there is a marked progression in the average (B − V)_{o} and (V − R)_{o} colors of RSGs in these three galaxies; with the higher metallicity systems having a later average spectral type, consistent with previous findings by Elias, Frogel, & Humphreys for the Milky Way and Magellanic Clouds. More significantly, we find that there is a clear progression with metallicity in the relative number of the highest luminosity RSGs, a trend that is apparent both in absolute visual magnitude and in bolometric luminosity. Thus any use of RSGs as distance indicators requires correction for the metallicity of the parent galaxy. Our findings are in accord with the predictions of the “Conti scenario” in which higher metallicities result in higher mass-loss rates, resulting in a star of a given luminosity spending an increasing fraction of its He-burning lifetime as a WR star rather than as a RSG with increasing metallicity. The fact that the distribution of luminosities have extended tails extending to higher luminosity (with the possible exception of the M 31 sample, which contains only one high luminosity RSG), suggests that many WRs may pass through a RSG phase, albeit for a short period. We find that over the factor of 5 difference in metallicity from NGC 6822 to M 31 the masses of the most luminous RSGs change by about a factor of two, with the highest mass RSGs having masses of 25 − 30M_{⊙} in NGC 6822, 18M_{⊙} in M33, and 13−15M_{⊙} in M31. These masses are lower than the usually assumed limits for evolution to the WR stage, but are not necessarily in conflict with the number of unevolved O and B stars if the time spent as a He-burning object is split between the RSG and WR phases over a large mass range.

Accepted by: The Astrophysical Journal For preprints, contact massey@noao.edu

Also available from the URL ftp://tofu.tuc.noao.edu//pub/rsgrevpre.ps
or by anonymous ftp at tofu.tuc.noao.edu, cd pub, get rsgrevpre.ps

Cyclicities in the light variations of LBVs. II. R 40 developing an S Dor phase

C. Sterken^{1}, M. de Groot^{2} and A.M. van Genderen^{3}

1 University of Brussels (VUB), Pleinlaan 2, 1050 Brussels, Belgium 2 Armagh Observatory, College Hill, Armagh BT61 9DG, Northern Ireland 3 Leiden Observatory, Postbus 9513, 2300RA Leiden, The Netherlands

Strömgren differential photometry of R 40 collected during the time interval 1986–1996 is analysed together with Walraven photometry. The gradual brightening of the star over the last 10 years can be described by a linear trend with superimposed oscillations (in v, b and y) with frequency 0.0008 cd^{−1} (∼ 1300 d cycle). We interpret these oscillations as “normal S Dor” phases, and suggest that the quasi-linear brightening of the star is the ascending branch of a growing very-long-term S Dor phase (VLT–SD), as found by van Genderen et al. (1997a) in AG Car and S Dor itself. As R 40 is now becoming fainter and bluer, the length of the VLT–SD cycle is about 20 years.

Accepted by: Astronomy & Astrophysics For preprints, contact csterken@vub.ac.be

Quasi-Periodic Occultation by a Precessing Accretion Disk and Other Variabilities of SMC X-1

P. S. Wojdowski^{1}, G. W. Clark^{1}, A. M. Levine^{1}, J. W. Woo^{2}, S. N. Zhang^{3}

1 Massachusetts Institute of Technology, USA 2 Harvard-Smithsonian Center for Astrophysics, USA 3 USRA/NASA-MSFC, USA

We have investigated the variability of the binary X-ray pulsar, SMC X-1, in data from several X-ray observatories. We confirm the 60-day cyclic variation of the X-ray flux in the long-term monitoring data from the RXTE and CGRO observatories. X-ray light curves and spectra from the ROSAT, Ginga, and ASCA observatories show that the uneclipsed flux varies by as much as a factor of twenty between a high-flux state when 0.71 second pulses are present and a low-flux state when pulses are absent. In contrast, during eclipses when the X-rays consist of radiation scattered from circumsource matter, the fluxes and spectra in the high and low states are approximately the same. These observations prove that the low state of SMC X-1 is not caused by a reduction in the intrinsic luminosity of the source, or a spectral redistribution thereof, but rather by a quasi-periodic blockage of the line of sight, most likely by a precessing tilted accretion disk. In each of two observations in the midst of low states a brief increase in the X-ray flux and reappearance of 0.71 second pulses occurred near orbital phase 0.2. These brief increases result from an opening of the line of sight to the pulsar that may be caused by wobble in the precessing accretion disk. The records of spin up of the neutron star and decay of the binary orbit are extended during 1991–1996 by pulse-timing analysis of ROSAT, ASCA, and RXTE PCA data. The pulse profiles in various energy ranges from 0.1 to > 21 keV are well represented as a combination of a pencil beam and a fan beam. Finally, there is a marked difference between the power spectra of random fluctuations in the high-state data from the RXTE PCA below and above 3.4 keV. Deviation from the fitted power law around 0.06 Hz may be QPO.

Accepted by: The Astrophysical Journal For preprints, contact pswoj@space.mit.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9801168
or by anonymous ftp at ftp://space.mit.edu/pub/pswoj/smcx1/disk.ps.gz

Ca II Triplet Spectroscopy of Giants in SMC Star Clusters: Abundances, Velocities and the Age-Metallicity Relation

G. S. Da Costa^{1} and D. Hatzidimitriou^{2}

1 Mount Stromlo and Siding Spring Observatories, The Australian National University, Private Bag, Weston Creek Post Office, ACT 2611, Australia 2 Physics Department, University of Crete, P.O. Box 2208, Heraklion 710 03, Crete, Greece

We have obtained spectra at the Ca II triplet of individual red giants in seven SMC star clusters whose ages range from ∼4 to 12 Gyr. The spectra have been used to determine mean abundances for six of the star clusters to a typical precision of 0.12 dex. When combined with existing data for other objects, the resulting SMC age-metallicity relation is generally consistent with that for a simple model of chemical evolution, scaled to the present-day SMC mean abundance and gas mass fraction. Two of the clusters (Lindsay 113 and NGC 339), however, have abundances that are ∼0.5 dex lower than that expected from the mean age-metallicity relation. It is suggested that the formation of these clusters, which have ages of ∼5 Gyr, may have involved the infall of unenriched gas, perhaps from the Magellanic Stream. The spectra also yield radial velocities for the seven clusters. The resulting velocity dispersion is 16 ± 4 km s^{−1}, consistent with those of the SMC planetary nebula and carbon star populations.

Accepted by: The Astronomical Journal For preprints, contact gdc@mso.anu.edu.au

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9802008
or by anonymous ftp at mso.anu.edu.au/pub/gdc/docs/smc amr.ps.gz

The evolution of the V − K colours of single stellar populations

Leo Girardi^{1} and Gianpaolo Bertelli^{2}

1 Max-Planck-Institut für Astrophysik, Karl-Schwarzschild-Str. 1, D-87540 Garching, Germany 2 CNR, and Department of Astronomy, Vicolo dell’Osservatorio 5, I-35122 Padova, Italy

We discuss how the integrated V − K colours of single stellar populations evolve with age and metallicity. Some of the uncertainties associated to the properties of the underlying stellar models are thoroughly discussed. First, we present the effects that different formulations for the mass-loss rates produce on the initial-final mass relation, and hence on the integrated colours. Second, we describe the colour changes occurring at the onset of the AGB and RGB stars. The onset of the AGB is marked by a colour jump to the red, whose amplitude is highly dependent on the metallicity and mass-loss rates adopted in the models. The overluminosity with respect to the standard core mass–luminosity relation, that occurs in the most massive AGB stars, causes a smoothing of this colour jump. On the other hand, we find that a transient red phase takes place sim1.5×10^{8} yr after the RGB develops. We then discuss the possible reasons for the transition of V − K colours (from ∼ 1.5 to 3) that takes place in LMC clusters of SWB type IV. A revision of the ages attributed to the clusters reveals that the transition may not be as fast as originally suggested. A gradual (or delayed) transition of the colours, as predicted by models which include the overluminosity of the most massive AGB stars, describes better the data than the sudden colour-jump predicted by classical models.

Submitted to: Monthly Notices of the Royal Astronomical Society

For preprints, contact leo@mpa-garching.mpg.de
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9801145

Chemical Evolution of the Magellanic Clouds: analytical models

B.E.J. Pagel^{1} and G. Tautvaišienė^{2}

1 NORDITA, Blegdamsvej 17, Dk-2100 Copenhagen, Denmark 2 Institute of Theoretical Physics and Astronomy, Goštauto 12, Vilnius 2600, Lithuania

We have extended our analytical chemical evolution modelling ideas for the Galaxy (Pagel & Tautvaišienė 1995, 1997) to the Magellanic Clouds. Unlike previous authors (Russell & Dopita 1992; Tsujimoto et al. 1995; Pilyugin 1996), we assume neither a steepened IMF nor selective galactic winds, since among the α-particle elements only oxygen shows a large deficit relative to iron and a similar deficit is also found in Galactic supergiants. Thus we assume yields and time delays identical to those that we previously assumed for the solar neighbourhood. We include inflow and non-selective galactic winds and consider both smooth and bursting star formation rates, the latter giving a better fit to the age-metallicity relations. We predict essentially solar abundance ratios for primary elements and these seem to fit most of the data within their substantial scatter. Our LMC model also gives a remarkably good fit to the anomalous Galactic halo stars discovered by Nissen & Schuster (1997).

Our models predict current ratios of SNIa to core-collapse supernova rates enhanced by 50 per cent and 25 per cent respectively relative to the solar neighbourhood, in fair agreement with ratios found by Cappellaro et al. (1993) for Sdm-Im relative to Sbc galaxies, but these ratios are sensitive to detailed assumptions about the bursts and a still higher enhancement in the LMC has been deduced from X-ray studies of remnants by Hughes et al. (1995). The corresponding ratios integrated over time up to the present are slightly below 1, but they exceed 1 if one compares the Clouds with the Galaxy at times when it had the corresponding metallicities.

Submitted to: Monthly Notices of the Royal Astronomical Society For preprints, contact pagel@nordita.dk

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9801221

Thesis Abstract

Development of the Eros-2 detector for microlensing searches

Comparison of LMC and SMC cepheids

Florian Bauer^{1,2}

CEA, DSM, DAPNIA, Centre d’Études de Saclay, F-91191 Gif-sur-Yvette Cedex, France 2 Collège de France, LPC, IN2P3-CNRS, 11 place Marcellin Berthelot, 75231 Paris Cedex, France

The EROS collaboration has been involved since 1990 in the search for dark baryonic objects in the halo of the Milky Way using microlensing of stars in the Magellanic Clouds. In 1994 the Eros collaboration built up a new detector, in order to increase the sensitivity to microlensing events by more than one order of magnitude.

The new astronomical wide field imager consists of two 4k × 8k CCD mosaic cameras, which are mounted on a 1 m Ritchey–Chretien telescope. The optical mounting allows simultaneous imaging in two focal planes with different color passbands.

A modular electronic system permits an individual computer controlled setting and a parallel readout of the CCDs. The readout noise is around 6-7 e^{−}, the cross-talk better than 10^{−5}, the readout time is 50 seconds for both mosaics. The whole Eros-2 system saw first light on June 1996.

The survey of millions of stars every night allows to search for variable stars, such as cepheids. After a massive photometric reduction and a periodicity search, 300 cepheids were found towards the LMC and 620 towards the SMC over two 1.5 square degree fields. These catalogs allow to undertake a comparison between the cepheids in both clouds; they are complete for periods ranging between 0.5 and 30 days.

For the first time a change of the slope in the PLC “plane” has been observed for fundamental mode cepheids with periods smaller than 2.5 days toward the SMC. No change in the slope could be observed for LMC cepheids. Different experimental biases were considered as a possible origin for this effect; they could all be discarded. This phenomenon could be due to evolutionary effects combined with convection.

Furthermore different resonances between pulsation modes have been observed, three new double mode cepheids have been found and two cepheid-rich clusters have been discovered, which contain eight cepheids each. Finally a study of the scatter of the period–luminosity relation is presented, which permits to estimate the LMC depth dispersion.

PhD thesis (in French), Université Paris 7, France defended on December 9th, 1997 Supervisor: Pierre Bareyre

For copies or related preprints, contact florian@hep.saclay.cea.fr

Meeting Announcement

ESO Conference on: Chemical Evolution from Zero to High Redshift

Scope, date and venue A 3-day conference on observational determination and interpretation of chemical abundances and evolution from stars, interstellar medium and local group galaxies at zero redshift to distant galaxies, clusters of galaxies and intergalactic medium at high redshift is to be held in Garching, Germany on 14-16 October 1998.

The broad aim of the conference is to assess the interplay between methods of measuring chemical abundances with the astrophysical models of galaxy evolution. Particular emphasis is to be given to discussion of the intrinsic uncertainties and merits of the various abundance determination methods, as well as on the prospects for future instrumentation.

Outline of the scientific program Anticipated topics:

• Solar and Solar System abundances

• Agreement between Solar Neighbourhood stellar and ISM abundances

• Abundance patterns in the Galaxy and comparison of observations with models

• Abundances of integrated populations in the Local Group

• Abundance fluctuations and gradients in galaxies

• Abundances anomalies in the nearby Universe

• Trends in abundances in clusters of galaxies

• Abundances from QSO absorption lines

• Abundances in high z starburst galaxies

• Dust and molecules at high z

• New windows - NGST, FIRST, MMA, etc.

Scientific Organizing Committee F. Combes (Obs. de Paris, Meudon); D. Garnett (Univ. Minnesota, Minneapolis); G. Kauffmann (MPA, Garching); C. Leitherer (STScI, Baltimore); D. Lennon (Univ. Sternwarte, Munich); J. Mathis [Chair] (Univ. Wisconsin, Madison); M. Pettini (RGO, Cambridge); M. Rosa (ST-ECF, Garching); P. Shaver (ESO, Garching); E. Terlevich (RGO, Cambridge) D. Tytler (Univ. California, San Diego)

Web page Further information can be found on the conference Web page:

http://www.eso.org/gen-fac/meetings/chemi98/ .

Please contact Gertrud Contardo (gcontard@eso.org) for details and registration and John Mathis

(mathis@madraf.astro.wisc.edu) for scientific input.

Local Organizing Committee G. Contardo, C. Stoffer, J. R. Walsh (all ESO, Garching)